# KAN-ODEs: Kolmogorov-Arnold Network Ordinary Differential Equations for Learning Dynamical Systems and Hidden Physics

### v2

@@ I Introduction
In contrast to Neural ODEs, physics-informed neural networks (PINNs) leverage the governing equations of the dynamical system, wherever available, in the loss function to train a traditional neural network that maps the time and spatial coordinates directly to the state variables [6, 12, 13]. This physics-embedded loss function allows for training that aims to fit both the data and the known governing equations without having to directly solve what are typically expensive PDEs. These networks are significantly interpretable thanks to the governing equations used as a soft constraint in the training cycle, giving physical meaning to trends and patterns that emerge from the training process. The inclusion of the governing equations as a soft constraint also allows the PINN to train with relatively small datasets, which is a useful benefit in many scientific and engineering applications with high costs of data acquisition [6]. Other frameworks have also been developed using a physics-based structure similar to PINN. The auto-regressive dense encoder-decoder model [14] encodes physics into a time-integration network scheme, while the physics-informed convolutional-recurrent network [15] runs PDE solutions through convolutional filters with physical penalty terms and hard-encoded boundary and initial conditions to efficiently develop models. Other works have also investigated the use of hard constraints for PINN-type models for initial and boundary conditions with strong results [16, 17]. PINN-type loss functions have also been included in Neural ODE frameworks [18]. Further combinations also exist in the realm of physics-based Neural ODEs, such as Chemical Reaction Neural Networks [19], which hard constrain Neural ODE frameworks to directly learn governing equation parameters from real [20, 21] and noisy [22, 23] experimental data. A downside of these physics-based techniques is their requirement of strong prior knowledge of the system’s governing laws. In cases where such knowledge is not available, other methods are required.
@@
In this study, we introduce the possibility of learning interpretable and modular ODE and PDE models for dynamical systems without prior knowledge or functional form assumptions by leveraging KANs as gradient-getters within the concept of Neural ODEs. In doing so, we propose a combined framework that benefits from the high-accuracy, parameter-learn, and strongly interpretable KAN network structure coupled with the general-purpose, grid-independent, and solver-flexible Neural ODE framework. A qualitative depiction of the comparison between the currently proposed KAN-ODE method and the other discussed data-driven approaches for dynamical system modeling is shown in Fig. 1. While the current state of the art is generally situated on the diagonal of this plot, where increased prior knowledge fed into a machine learning method delivers better interpretability, we place KAN-ODEs slightly below the diagonal, thanks to their potential to deliver interpretable activations and even symbolic relationships without requiring prior knowledge or presumed functional forms in the training cycle. The ODE-based formulation is also inherently flexible, allowing for variable timesteps depending on the gradient landscape as well as the application of the KAN-ODE as a component of a larger-scale model, for example as the source term in a nonlinear transport PDE. We demonstrated the promise of KAN-ODEs through a comprehensive comparison against standard Neural ODEs, as well as an array of tests at different scales and levels of complexity including a predator-prey ODE, wave propagation and shock formation PDEs, the complex Schrödinger equation, and a phase separation PDE. We found that KAN-ODEs succeed in all of these cases and are able to reconstruct complete temporal profiles even when given sparse training data. KAN-ODEs with a small number of parameters were able to converge with remarkable accuracy, and sparsity or symbolic relationships imposed during the training cycle or during postprocessing were able to give KAN-ODEs substantial generalization capability. Overall, we found that this encoding of KANs into the Neural ODE framework is a promising step toward small-data and low-knowledge yet interpretable machine learning for dynamical system modeling.
@@ II.1 Kolmogorov Arnold Networks as Gradient Evaluators
Expanded mathematical formulations for KANs with multivariate outputs become increasingly bulky and cumbersome and are available in the extensive [26]. For convenience, we use the notation [$n_{l},n_{l+1},N$] to represent an $l$-th KAN layer for the remainder of this work, where $n_{l}$ is the input dimension, $n_{l+1}$ is the output dimension, and $N$ is the grid size. For example, a 2-input and 1-output KAN with a 10-node hidden layer and a 5-point grid is represented as ”[2, 10, 5], [10, 1, 5]”, for the connections between the input and hidden layer, and the hidden and output layer, respectively.
@@ II.2 KAN-Ordinary Differential Equations
where $\bm{\theta}$ is the collection of all trainable weights in the KAN. Here, the input and output dimensions of the KAN are fixed as the dimension of the state variables $\mathbf{u}$, and it is used as a gradient getter. Equation  7 can be solved using a traditional ODE solver from an initial time $t_{0}$ to a desired time $t$ such that
@@
where $\mathbf{u}^{\text{KAN}}$ is the prediction, $\mathbf{u}^{\text{obs}}$ is the observation, and $N$ is the total number of time steps. To optimize $\mathcal{L}$, we require gradients with respect to $\bm{\theta}$. Generally, two options are available with the forward sensitivity method [8, 34] and the adjoint sensitivity method [7, 35]. In this paper, the adjoint sensitivity method is adopted to efficiently handle a potentially large Kolmogorov-Arnold network [35, 8, 7] thanks to its more advantageous scaling with model size. By introducing an adjoint state variable $\bm{\omega}$, the augmented system equation can be derived, which allows for gradient backpropagation through the ODE integrators.
@@
where $\mathbf{z}(t)=\mathbf{u}(-t)$, $\mathbf{g}=\text{KAN}$, and $T$ denotes the transpose. After Eq. 10 is computed, $\bm{\theta}$ is updated using a gradient descent method. A schematic depicting the overall training cycle is provided in Fig. 2.
@@ III Experiments
We present five inference examples in this work to demonstrate the broad capability of KAN-ODEs, grouped into three subsections. We first validated our methodology and compared its performance against the standard MLP-based Neural ODE in Sec. III.1, where the dynamics of the Lotka-Volterra predator-prey model are inferred. There we also comprehensively explored the neural scaling benefits of KAN-ODEs, as well as the efficient representations that can be inferred via sparsification and pruning. Then, we tackled a one-dimensional PDE in Sec. III.2, using a KAN-ODE to infer a symbolic source term in a wave propagation simulation. This example demonstrates the flexibility of KAN-ODEs to be used as submodels in higher complexity simulations, as well as their capability to directly extract symbolic models from data. Finally, we demonstrated the scalability and inference power of KAN-ODEs as standalone models in Sec. III.3 by inferring complete data-driven solutions to PDEs using limited samples. Examples here include shock wave formation governed by the Burgers’ equation and quantum wave function evolution governed by the Schrödinger equation, with an additional phase separation case using the Allen-Cahn equation provided in Appendix C.
III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra
In this section, we targeted the Lotka-Volterra predator-prey model shown in Eq. 11. We replicated the simulation parameters used in the Neural ODE study of [9]. Namely, we used $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ with an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$ and a time span of $t\in[0,14]$ s. The first 3.5 s were used to train the KAN-ODE (Fig. 3(A)), while the remainder of the time history was withheld to validate the KAN-ODE’s ability to forecast the states at unseen times. The temporal grid used for training and loss evaluation had a spacing of 0.1 s. We tested the KAN-ODEs with various architectures shown in Table 1. The inputs to each KAN layer in this example were normalized to be on the $[-1,1]$ range required by the RBF networks via the hyperbolic tangent function, as in [32].
@@ III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra
Loss profiles from training a 240-parameter KAN-ODE of shape [2, 10, 5], [10, 2, 5] as a representative case are shown in Fig. 3(B1). Strong convergence is observed at or even before $10^{4}$ epochs into training, while convergence into the $10^{-7}$ range of MSE loss occurs toward $10^{5}$ epochs. The converged KAN-ODE prediction is shown in Fig. 3(A), where the agreement in the testing window ($t\in[3.5,14]$) is nearly indistinguishable from the agreement in the training window ($t\in[0,3.5]$). We remark again that the trained KAN-ODE is reconstructing both the seen training data and unseen testing data given only the initial condition ($\mathbf{u}_{0}=[1,1]$), and can deliver output values on any temporal grid, thanks to its coupling to an ODE solver.
@@ III.1.1 Benchmark Tests and Neural Scaling Behavior
In this 1-to-1 comparison using a data generation and training regime borrowed from a well-cited Neural ODE work, our proposed KAN-ODE beats the performance of the Neural ODE in all key metrics. Inspection of the loss profiles of Figs. 3(B1-B2) shows that the KAN-ODE converged to a substantially smaller training loss of $8.3\times 10^{-7}$ over $10^{5}$ epochs compared to the similarly-sized Neural ODE, which reached only $3\times 10^{-5}$ in the same number of epochs. The KAN-ODE qualitatively appears to overfit more than the Neural ODE model, especially in the later epochs. Quantitatively, however, it still beats the Neural ODE in accuracy to the unseen testing data. Even in the last 10% of the training cycle, i.e. epochs $90,000-100,000$, where the KAN-ODE appeared to overfit substantially, it still achieved an average testing loss of $6.8\times 10^{-3}$, with a minimum value in that range of $1.9\times 10^{-5}$. The Neural ODE in this same range averaged a much larger $1.4\times 10^{-2}$, with a minimum value of $5.4\times 10^{-5}$. If training for the KAN-ODE was stopped early at $2\times 10^{4}$ epochs when the testing error just begins to creep upward, these statistics would tend even more strongly in its favor.
@@
Figure 4: Sparsification, pruning, symbolic regression, and generalization results for Lotka-Volterra dynamics. (A) KAN with 72 parameters ([2, 3, 5], [3, 2, 5]), pruned from an initial 240 parameters ([2, 10, 5], [10, 2, 5]). (B) Symbolic KAN derived from (A), where each of the twelve activations is replaced by a univariate symbolic expression. (C) Generalization error when extrapolating outside of the yellow $(x,y)$ contour explored in the training and testing data. Each column represents a different gradient getter and has one row each for $dx/dt$ and $dy/dt$. Generalization is seen to improve when moving from MLP gradients to increasingly sparse KANs, and finally to symbolic representations fitted to KANs.
@@ III.1.2 Interpretation and Generalization of KAN-ODEs with Varying Sizes
In this section, we implemented sparsification, pruning, and symbolic regression to demonstrate a framework for developing interpretable KAN-ODE solvers. We additionally heuristically explored the generalizability of the various MLP and KAN-based models, where we found that KAN-ODEs have strong generalization potential even with limited training data. In addition to the 252-parameter Neural ODE and 240-parameter KAN-ODE as reference cases, three models were studied: a sparse KAN-ODE trained via parameter regularization and pruning, a symbolic KAN developed via symbolic regression on each univariate activation of the sparse KAN-ODE, and a symbolic representation for the entire model via symbolic regression on the global input-output pairs of the sparse KAN-ODE.
@@
which correspond nearly identically to our model parameters of 1.5, 1, 1, and 3. This result is effectively the pinnacle of interpretability and accuracy, though we believe is not a realistic option for most systems of larger-scale merit. Complete symbolic regression on an entire learned model or dataset is often not feasible, especially for complex models with higher dimensional inputs and outputs than are studied in this relatively straightforward ODE system [26]. Even when numerically manageable, clean convergence to compact functional forms is not guaranteed. On the other hand, a symbolic KAN like that shown in Fig. 4(B) requires symbolic regression only on a series of straightforward univariate activations and retains entirely interpretable intermediate steps. We present this global symbolic representation result as an interesting aside and a benchmark for comparison, but propose that sparse KAN-ODEs and symbolic KAN-ODEs are the most likely to have broader applicability in more realistic systems depending on the desired degree of interpretability and accuracy.
@@
We move next to a discussion of the generalizability of KAN-ODEs. To conform to prior studies and faciliate fair comparison against MLP-based Neural ODEs, all models in this section were trained on data generated from a single initial condition, and tested on a continuation of these predator-prey oscillations for three more complete cycles (see Fig. 3). There exist many other potential $(x,y)$ combinations, however, and here we heuristically tested these trained models on their ability to generalize to completely unseen conditions.
@@
The gradient error landscapes of the five models studied here (MLP-NODE, Dense KAN-ODE, Sparse KAN-ODE, Symbolic KAN-ODE, and symbolic representation of KAN-ODE) are shown from left to right, respectively, in Fig. 4(C), where the top row is the $dx/dt$ error and the bottom row is the $dy/dt$ error. The MLP-NODE results in the first column show near-zero error in the immediate vicinity of the training data contour, as expected from the performance metrics of Fig. 3. However, they also show significant gradient error in almost all ($x,y$) ranges not directly on or immediately near the training data contour, revealing the expected weak generalizability of the standard Neural ODE due to the fact that training data is limited to just those points on this contour.
@@
The gradient error landscape of the dense KAN-ODE beats the MLP in generalizability, with much smaller gradient error values throughout the domain and especially inside of the yellow training contour where not only is the overall error substantially reduced, but we begin to additionally observe the formation of thin bands of near-perfect reconstruction away from the data contour. This result is rather interesting, as it was originally proposed in [26] that KANs beat MLPs in generalization thanks to their smaller number of parameters, but we observe here a substantial improvement with similarly sized models. In any case, these results suggest the potential of KAN-ODEs in the development of powerful models from limited data.
@@
The gradient error landscape of the sparse KAN-ODE shows improvements over the dense KAN-ODE landscape. Inside the yellow data contour, we see further formation of near-perfect gradient reconstruction bands, surrounded by relatively low error regions. Results toward the edges of the landscape are mixed: the $dx/dt$ reconstruction captures the top left of the ($x,y$) domain with excellent accuracy but suffers from a large error in the upper right portion of the domain, while the $dy/dt$ reconstruction performs well toward low $y$ values, but suffers at high $y$ values. On the whole, we remark an overall improvement in generalization despite a nearly 4x reduction in KAN size, indicating the utility of sparse KAN-ODEs not just for compact and interpretable models, but also for improved generalization and extrapolation capabilities thanks to their smaller size. This result corroborates similar observations from [26] of the increase in generalizability with decreased KAN parameter counts.
@@
The Symbolic KAN-ODE sees remarkable improvement in the $dx/dt$ generalization, with the low-error area being pushed out significantly toward high $x$ and $y$ values, further growth of the white bands of near-zero error throughout the domain, and a general lightening of the plot indicating better extrapolation and generalization. The $dy/dt$ reconstruction sees modest improvement, especially inside the data contour. We propose that this generalization improvement occurs due again to the more compact forms present in the symbolic KAN-ODE when compared to even the sparse KAN-ODE, and the extrapolation benefits that come with smaller system representations [26]. In the significantly out-of-training regions of $dx/dt$ where the symbolic KAN-ODE succeeds and the gridded basis functions of the dense KAN-ODE and Sparse KAN-ODE fail, we suggest that the input samples may be exiting the trained region of the RBF basis functions yet remaining within the useful range of the expressions used in the symbolic KAN-ODE.
@@
The symbolic representation of the entire KAN obviously generalizes with near-perfect accuracy thanks to its discovery of what is essentially the complete underlying model. This result is plotted in the final column of Fig. 4(C) for reference, but we again emphasize that the other three KAN models are more reasonable choices for realistic applications: Dense KAN-ODEs succeeded with remarkably low training and testing error, sparse KAN-ODEs retained much of this low error while appearing to improve generalizability thanks to their smaller size, and Symbolic KAN-ODEs were substantially more interpretable and human-readable while also appearing to improve generalizability thanks to their smaller size and replacement of limited-scope basis functions with standard symbolic functions.
@@
In this Lotka-Volterra example, we systematically and comprehensively explored the use of KAN-ODEs in a straightforward ODE system and benchmarked their performance against comparable MLP-based Neural ODEs in terms of accuracy, speed, model size, convergence rate, interpretability, and generalization. We found that KAN-ODEs are an extremely efficient method to represent the predator-prey dynamics in this case, and beat Neural ODEs in all relevant quantitative metrics. We additionally explored the interpretability of KAN-ODEs via symbolic regression of each activation function, and then heuristically evaluated a series of MLP and KAN-based ODE approaches in terms of generalizability to unknown data, where we again found KAN-ODEs superior. In the next sections, we present larger-scale examples to showcase the flexibility and inference capabilities of KAN-ODEs.

### v3

@@ I Introduction
In contrast to Neural ODEs, physics-informed neural networks (PINNs) leverage the governing equations of the dynamical system, wherever available, in the loss function to train a traditional (i.e. not ODE based) neural network that maps the time and spatial coordinates directly to the state variables [6, 12, 13]. This physics-embedded loss function allows for training that aims to fit both the data and the known governing equations without having to directly solve what are typically expensive PDEs. These networks are significantly interpretable thanks to the governing equations used as a soft constraint in the training cycle, giving physical meaning to trends and patterns that emerge from the training process. The inclusion of the governing equations as a soft constraint also allows PINNs to train with relatively small datasets, which is a useful benefit in many scientific and engineering applications with high costs of data acquisition [6]. Other frameworks have also been developed using a physics-based structure similar to PINN. The auto-regressive dense encoder-decoder model [14] encodes physics into a time-integration network scheme, while the physics-informed convolutional-recurrent network [15] runs PDE solutions through convolutional filters with physical penalty terms and hard-encoded initial and boundary conditions to efficiently develop models. Other works have also investigated the use of hard constraints with PINN-type models for initial and boundary conditions with strong results [16, 17]. PINN-type loss functions have additionally been included in Neural ODE frameworks [18]. Further combinations exist in the area of physics-based Neural ODEs, such as Chemical Reaction Neural Networks [19], which hard constrain Neural ODE frameworks to directly learn governing equation parameters from real [20, 21] and noisy [22, 23] experimental data. A downside shared throughout these numerous physics-based techniques is their requirement of strong prior knowledge of the system’s governing laws. In cases where such knowledge is not available, other methods are required.
@@
In this study, we introduce the possibility of learning interpretable and modular ODE and PDE models for dynamical systems without prior knowledge or functional form assumptions by leveraging KANs as gradient-getters within the concept of Neural ODEs. In doing so, we propose a combined framework that benefits from the high-accuracy, parameter-lean, and strongly interpretable KAN network structure coupled with the general-purpose, grid-independent, and solver-flexible Neural ODE framework. A qualitative depiction of the comparison between the currently proposed KAN-ODE method and the other discussed data-driven approaches for dynamical system modeling is shown in Fig. 1. While the current state of the art is generally situated on the diagonal of this plot, where increased prior knowledge fed into a machine learning method delivers better interpretability, we place KAN-ODEs slightly below the diagonal, thanks to their potential to deliver interpretable activations and even symbolic relationships without requiring prior knowledge or presumed functional forms in the training cycle. The ODE-based formulation is also inherently flexible, allowing for variable timesteps depending on the gradient landscape as well as the application of the KAN-ODE as a component of a larger-scale model, for example as the source term in a nonlinear transport PDE. We demonstrated the promise of KAN-ODEs through a comprehensive comparison against standard Neural ODEs, as well as an array of tests at different scales and levels of complexity including a predator-prey ODE, wave propagation and shock formation PDEs, the complex Schrödinger equation, and a phase separation PDE. We found that KAN-ODEs succeed in all of these cases and are able to reconstruct complete temporal profiles even when given sparse training data. KAN-ODEs with a small number of parameters were able to converge with remarkable accuracy, and sparsity or symbolic relationships imposed during the training cycle or during postprocessing were able to give KAN-ODEs substantial generalization capability. Overall, we found that this encoding of KANs into the Neural ODE framework is a promising step toward small-data and low-knowledge yet interpretable machine learning for dynamical system modeling.
@@ II.1 Kolmogorov Arnold Networks as Gradient Evaluators
Expanded mathematical formulations for KANs with multivariate outputs become increasingly bulky and cumbersome and are available in the extensive [26]. For convenience, we use the notation [$n_{l},n_{l+1},N$] to represent an $l$-th KAN layer for the remainder of this work, where $n_{l}$ is the input dimension, $n_{l+1}$ is the output dimension, and $N$ is the grid size. For example, a 2-input and 1-output KAN with a 10-node hidden layer and a 5-point grid is represented as ”[2, 10, 5], [10, 1, 5]” for the connections between the input and hidden layer, and the hidden and output layer, respectively.
@@ II.2 KAN-Ordinary Differential Equations
where $\bm{\theta}$ is the collection of all trainable weights in the KAN. Here, the input and output dimensions of the KAN are fixed as the dimension of the state variables $\mathbf{u}$, and it is used as a gradient getter. Equation 7 can be solved using a traditional ODE solver from an initial time $t_{0}$ to a desired time $t$ such that
@@
where $\mathbf{u}^{\text{KAN}}$ is the prediction, $\mathbf{u}^{\text{obs}}$ is the observation, and $N$ is the total number of time steps. To optimize $\mathcal{L}$, we require gradients with respect to $\bm{\theta}$. Generally, two options are available: the forward sensitivity method [8, 34] and the adjoint sensitivity method [7, 35]. In this paper, the adjoint sensitivity method is adopted to efficiently handle a potentially large Kolmogorov-Arnold network [35, 8, 7] thanks to its more advantageous scaling with model size. By introducing an adjoint state variable $\bm{\omega}$, the augmented system equation can be derived, which allows for gradient backpropagation through the ODE integrators.
@@
where $\mathbf{z}(t)=\mathbf{u}(-t)$, $\mathbf{g}=\text{KAN}$, and $T$ denotes the transpose. Detailed derivations of the adjoint equations (Eq. 10) can be found in [36, 37, 38, 7]. After Eq. 10 is computed, $\bm{\theta}$ is updated using a gradient descent method. A schematic depicting the overall training cycle is provided in Fig. 2.
@@ III Experiments
We present five inference examples in this work to demonstrate the broad capability of KAN-ODEs, grouped into three subsections. We first validated our methodology and compared its performance against the standard MLP-based Neural ODE in Sec. III.1, where the dynamics of the Lotka-Volterra predator-prey model were inferred. There we also comprehensively explored the neural scaling benefits of KAN-ODEs, as well as the efficient representations that can be inferred via sparsification and pruning. Then, we tackled a one-dimensional PDE in Sec. III.2, using a KAN-ODE to infer a symbolic source term in a wave propagation simulation. This example demonstrates the flexibility of KAN-ODEs to be used as submodels in higher complexity simulations, as well as their capability to directly extract symbolic models from data. Finally, we demonstrated the scalability and inference power of KAN-ODEs as standalone models in Sec. III.3 by inferring complete data-driven solutions to PDEs using limited samples. Examples there include shock wave formation governed by the Burgers’ equation and quantum wave function evolution governed by the Schrödinger equation, with an additional phase separation case using the Allen-Cahn equation provided in Appendix C.
III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra Equations
In this section, we targeted the Lotka-Volterra predator-prey model shown in Eq. 11. We replicated the simulation parameters used in the Neural ODE study of [9]. Namely, we used $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ with an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$ and a time span of $t\in[0,14]$ s. The temporal grid used for data generation, training, and loss evaluation had a spacing of 0.1 s. The Tsit5 ODE integrator was used to solve the governing equations for the true ODE trajectory (i.e. the ”ground truth” data used to train and validate the KAN-ODE model). The first 3.5 s of this ground truth trajectory was used to train the KAN-ODE (Fig. 3(A)) via Eqs. 9 and 10, while the remainder of the time history (3.5 s to 14 s) was withheld to validate the KAN-ODE’s ability to forecast the states at unseen times. While the training data is reported at 0.1 s increments, the KAN-ODE’s use of the Tsit5 adaptive solver requires much finer fidelity in its gradient accuracy. We tested the KAN-ODEs with various architectures shown in Table 1. The inputs to each KAN layer in this example were normalized to be on the $[-1,1]$ range required by the RBF networks via the hyperbolic tangent function, as in [32].
@@ III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra Equations
Loss profiles from training a 240-parameter KAN-ODE of shape [2, 10, 5], [10, 2, 5] as a representative case are shown in Fig. 3(B1). Strong convergence is observed at or even before $10^{4}$ epochs into training, while convergence into the $10^{-7}$ range of MSE loss occurs toward $10^{5}$ epochs. The converged KAN-ODE prediction is shown in Fig. 3(A), where the agreement in the testing window ($t\in[3.5,14]$) is nearly indistinguishable from the agreement in the training window ($t\in[0,3.5]$). We remark again that the trained KAN-ODE is reconstructing both the seen training data (i.e. the ground truth ODE solution from 0 s to 3.5 s) and unseen testing data (3.5 s to 14 s) given only the initial condition ($\mathbf{u}_{0}=[1,1]$), and can deliver output values on any temporal grid, thanks to its coupling to an ODE solver.
@@ III.1.1 Benchmark Tests and Neural Scaling Behavior
In this 1-to-1 comparison using a data generation and training regime borrowed from a well-cited Neural ODE work, our proposed KAN-ODE beats the performance of the Neural ODE in all key metrics. Inspection of the loss profiles of Figs. 3(B1-B2) shows that the KAN-ODE converged to a substantially smaller training loss of $8.3\times 10^{-7}$ over $10^{5}$ epochs compared to the similarly-sized Neural ODE, which reached only $3\times 10^{-5}$ in the same number of epochs. Both loss profiles exhibit significant oscillatory behavior while still successfully training, which we theorize may be due to the data-lean, large-batch regime studied here (i.e. a single dataset, all time steps of which are optimized in one batch). These oscillations, which do not appear to affect the final result, are smaller in the KAN-ODE training cycle than in the Neural ODE training cycle. The KAN-ODE qualitatively appears to overfit more than the Neural ODE model, especially in the later epochs. Quantitatively, however, it still beats the Neural ODE in accuracy to the unseen testing data, indicating an objective improvement in performance when switching from the Neural ODE to the KAN-ODE. Even in the last 10% of the training cycle (epochs $90,000-100,000$), where the KAN-ODE experiences a notable difference between training and testing losses, it still achieved an average testing loss of $6.8\times 10^{-3}$, with a minimum value in that range of $1.9\times 10^{-5}$. The Neural ODE in this same range averaged a much larger $1.4\times 10^{-2}$, with a minimum value of $5.4\times 10^{-5}$. If training for the KAN-ODE was stopped early at $2\times 10^{4}$ epochs when the testing error just begins to creep upward, these statistics would tend even more strongly in its favor. At this early stopping point, little to no indication of overfitting is observed when comparing its own training and testing losses, and all loss metrics still outperform those of the Neural ODE (even when the latter is trained to the full $1\times 10^{5}$ epochs). The KAN-ODE solution of Fig. 3(A) can additionally be generated at arbitrary time steps (e.g. 0.01 s rather than 0.1 s) with no detectable change to the testing results, indicating no overfitting to the specific training timesteps used here.
@@
Figure 4: Sparsification, pruning, symbolic regression, and generalization results for Lotka-Volterra dynamics. (A) Sparse KAN with 72 parameters ([2, 3, 5], [3, 2, 5]), pruned from an initial 240 parameters ([2, 10, 5], [10, 2, 5]). (B) Symbolic KAN derived from (A), where each of the twelve activations is replaced by a univariate symbolic expression. (C) Generalization error when extrapolating outside of the yellow $(x,y)$ points explored in the training data. Each column represents a different gradient getter and has one row each for $dx/dt$ and $dy/dt$. Generalization is seen to improve when moving from MLP gradients to increasingly sparse KANs, and finally to symbolic representations fitted to KANs.
@@ III.1.2 Interpretation and Generalization of KAN-ODEs with Varying Sizes
In this section we implemented sparsification, pruning, and symbolic regression to demonstrate a framework for developing interpretable KAN-ODE solvers. We additionally heuristically explored the generalizability of the various MLP and KAN-based models, where we found that KAN-ODEs have strong generalization potential even with limited training data. In addition to the 252-parameter Neural ODE and 240-parameter KAN-ODE (trained in Sec. III.1.1) as reference cases, three models were studied: a sparse KAN-ODE trained via parameter regularization and pruning, a symbolic KAN-ODE developed via symbolic regression on each univariate activation of the sparse KAN-ODE, and a symbolic representation for the entire model via symbolic regression on the global input-output pairs of the sparse KAN-ODE.
@@
which correspond nearly identically to our model parameters of 1.5, 1, 1, and 3. This near-perfect result is effectively the pinnacle of interpretability and accuracy, though we believe is not a realistic option for most systems of larger-scale merit. Complete symbolic regression on an entire learned model or dataset is often not feasible, especially for complex models with higher dimensional inputs and outputs than are studied in this relatively straightforward ODE system [26]. Even when numerically manageable, clean convergence to compact functional forms is not guaranteed. On the other hand, a symbolic KAN like that shown in Fig. 4(B) requires symbolic regression only on a series of straightforward univariate activations and retains entirely interpretable intermediate steps. We present this global symbolic representation result as an interesting aside and a benchmark for comparison, but propose that sparse KAN-ODEs and symbolic KAN-ODEs are the most likely to have broader applicability in more realistic systems depending on the desired degree of interpretability and accuracy.
@@
We move next to a discussion of the generalizability of KAN-ODEs. To conform to prior studies and faciliate fair comparison against MLP-based Neural ODEs, all models used previously in this Lotka-Volterra section were trained on data generated from a single initial condition, as was done in the Neural ODE study of [9]. There exist many other potential $(x,y)$ combinations, however, and here we heuristically tested these trained models on their ability to generalize to completely unseen conditions.
@@
The gradient error landscapes of the five models studied here (MLP-NODE, Dense KAN-ODE, Sparse KAN-ODE, Symbolic KAN-ODE, and symbolic representation of KAN-ODE) are shown from left to right, respectively, in Fig. 4(C), where the top row is the $dx/dt$ error and the bottom row is the $dy/dt$ error. The MLP-NODE results in the first column show near-zero error in the immediate vicinity of the training data, as expected from the performance metrics of Fig. 3. However, they also show significant gradient error in almost all ($x,y$) ranges not directly on or immediately near the training data points, revealing the expected weak generalizability of the standard Neural ODE due to the fact that training data is limited to the marked points only.
@@
The gradient error landscape of the dense KAN-ODE beats the MLP in generalizability, with much smaller gradient error values throughout the domain and especially inside of the yellow training points where not only is the overall error substantially reduced, but we begin to additionally observe the formation of thin bands of near-perfect reconstruction away from the data points. This result is rather interesting, as it was originally proposed in [26] that KANs beat MLPs in generalization thanks to their smaller number of parameters, but we observe here a substantial improvement with similarly sized models. In any case, these results suggest the potential of KAN-ODEs in the development of powerful models from limited data. In a brief aside, we also remark on the current result’s support of our previous argument (in Sec. III.1.1) that the fully-trained dense KAN-ODE does not overfit to the training data points, given its near-perfect interpolation between these points along the training profile. This result is unsurprising given the adaptive time stepping of the Tsit5 solver, but is nonetheless a strong result, especially when combined with the additional visible improvement in generalization outside of the training profile. If overfitting becomes substantial in larger-scale KAN-ODE applications, there are a handful of general neural network techniques and Neural-ODE adapted techniques that can be adapted to address this [46, 47, 48] including early stopping and network reduction (as proposed here) as well as data expansion, noise injection, and dropout.
@@
The gradient error landscape of the sparse KAN-ODE shows improvements over the dense KAN-ODE landscape. Inside the yellow data points, we see further formation of near-perfect gradient reconstruction bands, surrounded by relatively low error regions. Results toward the edges of the landscape are mixed: the $dx/dt$ reconstruction captures the top left of the ($x,y$) domain with excellent accuracy but suffers from a large error in the upper right portion of the domain, while the $dy/dt$ reconstruction performs well toward low $y$ values, but suffers at high $y$ values. On the whole, we remark an overall improvement in generalization with a nearly 4x reduction in KAN size, indicating the utility of sparse KAN-ODEs not just for compact and interpretable models, but also for improved generalization and extrapolation capabilities thanks to their smaller size. This result corroborates similar observations from [26] of the increase in generalizability with decreased KAN parameter counts.
@@
The Symbolic KAN-ODE sees remarkable improvement in the $dx/dt$ generalization, with the low-error area being pushed out significantly toward high $x$ and $y$ values, further growth of the white bands of near-zero error throughout the domain, and a general lightening of the plot indicating better extrapolation and generalization. The $dy/dt$ reconstruction sees modest improvement, especially inside the data points. We propose that this generalization improvement occurs due again to the more compact forms present in the symbolic KAN-ODE when compared to even the sparse KAN-ODE, and the extrapolation benefits that come with smaller system representations [26]. In the significantly out-of-training regions of $dx/dt$ where the symbolic KAN-ODE succeeds and the gridded basis functions of the dense KAN-ODE and Sparse KAN-ODE fail, we suggest that the input samples may be exiting the trained region of the RBF basis functions yet remaining within the useful range of the expressions used in the symbolic KAN-ODE.
@@
The symbolic representation of the entire KAN obviously generalizes with near-perfect accuracy thanks to its discovery of the exact underlying model form (with near-perfect parameters). This result is plotted in the final column of Fig. 4(C) for reference, but we again emphasize that the other three KAN models are more reasonable choices for realistic applications: Dense KAN-ODEs succeeded with remarkably low training and testing error, sparse KAN-ODEs retained much of this low error while appearing to improve generalizability thanks to their smaller size, and Symbolic KAN-ODEs were substantially more interpretable and human-readable while also appearing to improve generalizability thanks to their smaller size and replacement of limited-scope basis functions with standard symbolic functions.
@@
In this Lotka-Volterra example, we systematically and comprehensively explored the use of KAN-ODEs in a straightforward ODE system and benchmarked their performance against comparable MLP-based Neural ODEs in terms of accuracy, speed, model size, convergence rate, interpretability, and generalization. We found that KAN-ODEs are an extremely efficient method to represent the predator-prey dynamics of this case, and beat Neural ODEs in all relevant quantitative metrics. We additionally explored the interpretability of KAN-ODEs via symbolic regression of each activation function, and then heuristically evaluated a series of MLP and KAN-based ODE approaches in terms of generalizability to unknown data, where we again found KAN-ODEs superior. In the next sections, we present larger-scale examples to showcase the flexibility and inference capabilities of KAN-ODEs.